# Rocky planet formation in compact disks around M dwarfs

### v2

@@ Abstract
Context. Due to the improvements in radial velocity and transit techniques, we know that rocky or rocky/icy planets, in particular close-in super-Earths in compact configurations, are the most common ones around M dwarfs. On the other hand, thanks to the high angular resolution of ALMA we know that many disks around very low mass stars (between 0.1 and 0.5 M⊙) are rather compact and small (without observable substructures and radius less than 20 au), which favours the idea of an efficient radial drift that could enhance planet formation in the terrestrial zone.
Aims. We aim to investigate the potential formation paths of the observed close-in rocky exoplanet population around M dwarfs, specially close-in super-Earths, assuming that planet formation could take place in compact disks with an efficient dust radial drift.
Methods. We developed N-body simulations that include a sample of embryos growing by pebble accretion exposed to planet-disk interactions, star-planet tidal interactions and general relativistic corrections that include the evolution of the luminosity, radius and rotational period of the star. For a star of 0.1 M⊙ we considered different gas disk viscosity and initial embryo distributions. We also explore planet formation by pebble accretion around stars of 0.3 M⊙ and 0.5 M⊙. Lastly, for each stellar mass, we run simulations that include a sample of embryos growing by planetesimal accretion instead of pebble accretion.
Results. Our main result is that the sample of simulated planets that grow by pebble accretion in a gas disk with low viscosity () can reproduce the close-in low-mass exoplanet population around M dwarfs in terms of multiplicity, masses and semi-major axis. Furthermore, we found that a gas disk with high viscosity () and thus lower pebble accretion rates, can not reproduce the observed planet masses as no planet more massive than 0.5 M⊕ could be formed in our simulations. Also, we show that planetesimal accretion favours the formation of smaller planets than the ones formed by pebble accretion. Whether this planet population truly exist still remains unknown with the current instrumental sensitivity.
Conclusions. Rocky planet formation around M dwarfs can take place in compact and small dust disks driven by an efficient radial drift in a gas disk with low viscosity (). This result points towards a new approach in the direction of the disk conditions needed for rocky planet formation around very low mass stars.
@@
M dwarfs are the most abundant objects in our solar neighborhood, representing almost 70 of the stars in our Galaxy (Chabrier, 2003). Thanks to the improvements in radial velocity and transit detection techniques, we know that they are hosts to closely-packed planetary systems with multiple rocky planets (e.g. Shields et al., 2016; Mulders, 2018). Up to date, rocky planets are the most abundant around M dwarfs (Sabotta et al., 2021), frequently found close-in to the star (Ment & Charbonneau, 2023). Such population is of great interest as many of them are located into the habitable zone of their systems (e.g. Kasting et al., 1993; Kopparapu et al., 2013), which makes M dwarfs ideal targets for the search of life in the Universe.
@@
Even though several authors addressed the study of Earth-like planets around M dwarfs with both population synthesis model and N-body simulations (e.g. Dugaro et al., 2016; Coleman et al., 2019; Miguel et al., 2019; Liu et al., 2020; Sánchez et al., 2022), super-Earth (planets with sizes between Earth and Neptune) formation around M dwarfs is still under debate as it is challenging to form such planets in situ with the classic planet formation theory and based on the low disk masses typically derived from observations (e.g. Raymond et al., 2007; Ciesla et al., 2015; Schlichting, 2018). Thus, it is necessary to address the problem taking into account different premises. For example, Liu et al. (2019) developed a pebble-driven core accretion model in a population synthesis code to study the formation and evolution of planets around stars from the sub-stellar mass limit at 0.08 M⊙ up to 1 M⊙. They found that the characteristic mass of super-Earth is set by the pebble isolation mass, and that it increases linearly with the mass of the stellar host, and that planets with masses less than 20 M⊕ can be formed around stars with a wide range of metallicities. However, they did not consider gravitational interactions among the sample of embryos, which is of great importance as collisions among embryos play a primary role in defining the final masses and orbital configurations of planetary systems (e.g. Chambers, 2001; O’Brien et al., 2006; Raymond et al., 2007; de Elía et al., 2013; Ronco & de Elía, 2018; Sánchez et al., 2022). Thus, their simulations may not be realistic when compared to the final exoplanet properties. Other studies, such as Mulders et al. (2021), proposed a pebble-drift and accretion model to explain the occurrence rate of transiting planets around M dwarfs. They found that the fraction of close-in super-Earths is higher around lower-mass stars and matches the exoplanet occurrence rates from Kepler. But their results are based on the study of a two-planet system, and they do not take into account planet-disk interactions. The inclusion of gas disk interactions with a sample of embryos is crucial, as is defining the early dynamical evolution and orbital configuration of planetary systems (e.g. Tanaka & Ward, 2004; Paardekooper et al., 2011; Ida et al., 2020). Moreover, the inclusion of star-planet tidal interactions is also relevant in the long-term dynamical evolution of close-in planets around very low mass stars due to the fast evolution of the stellar radius and rotation period (e.g. Bolmont et al., 2011; Sánchez et al., 2020).
@@
Generally, the assumptions on the disk properties in planet population models may not be representative when compared with the latest insights from disk observations. Lately, the Atacama Large
@@
Millimeter/sub-millimeter Array (ALMA) disk surveys show the existence of compact dust disks around very low mass stars (e.g. Facchini et al., 2019; Kurtovic et al., 2021; van der Marel et al., 2022). A handful of disks are truly as small as 3-5 au in radius, whereas many of these disks remain unresolved in current ALMA images and their radius is only constrained to be 20 au. Such compact disks, lacking large scale substructures, are more often found around low-mass stars than high mass stars (van der Marel & Mulders, 2021) and are expected to be dominated by radial drift and thus a high pebble flux (Pinilla et al., 2013; Banzatti et al., 2020, 2023). In such disks, super-Earths are expected to form, consistent with a formation through pebble accretion in drift dominated disks (Mulders et al., 2021; van der Marel & Mulders, 2021). Recently, Chachan & Lee (2023) used a simple framework that combines the theory of pebble accretion with the measurements of dust masses in protoplanetary disks around late M dwarfs. They demonstrated that the
@@
FGK dwarfs (e.g. Dressing et al., 2015; Mulders et al., 2015), in agreement with (van der Marel & Mulders, 2021). Additionally, by modelling the evolution of a stellar cluster, Appelgren et al. (2023) investigated the efficiency of radial drift in clearing the disk of mm-sized particles. Along the disk lifetime, they found an agreement between the mass locked in pebbles and the evolution of dust masses and ages inferred from nearby star-forming regions. They stated that discs are born with dust masses up to around 50 M⊕ for low-mass stars of 0.1 M⊙ which quickly decreases to 1 M⊕ at 1-2 Myr due to radial drift, which suggests that at the earliest stages could be enough material to form inner planetary cores. Such high disk dust masses have indeed been inferred for embedded disks in the Class I stage (Tychoniec et al., 2020).
Motivated by the fact that the formation of planets more massive than the Earth is hard to explain with classical planet formation models, and that dust disks around very low mass stars are rather small and compact, dominated by radial drift and with an initial dust mass high enough to form inner planetary cores, we investigate different rocky planet formation paths around M dwarfs. In order to do that, we assume that planet formation takes place in compact dust disks, dominated by radial drift rather than the well-studied gapped disks (e.g. Andrews et al., 2018) where dust traps are thought to halt the radial drift (Pinilla et al., 2012). Based on this assumption, we developed different sets of N-body simulations that include a sample of small embryos growing by pebble accretion that experience planet-disk interactions as well as star-planet tidal interactions that incorporate the evolution of the luminosity, radius and rotational period of the star. All simulations run for 50 Myr to take into account late collisions among embryos, considering a gas disk lifetime of 10 Myr. We explore different planet formation scenarios around stars of 0.1 M⊙, 0.3 M⊙ and 0.5 M⊙.
This work is structured as follows. In Section 2 we describe the dust disk model used based on compact dust disk observations. In Section 3 we describe the planet formation model included in the N-body code and present different formation scenarios. Appendix A and B summarize the details on the gas disk evolution model and external forces on the N-body code. In Section 4 we show the resulting simulated planetary systems and explain their orbital history and final configurations. In Section 5 we compare the simulated planets with the confirmed close-in low-mass exoplanet population around M dwarfs. Finally, in Section 6 we summarize some points of discussion and in Section 7 we give the main conclusions of our work.
@@ 2 Compact dust disks: models and observations
Current ALMA observations of protoplanetary disks around M dwarfs at 1-2 Myr primarily show compact dust continuum disks which remain unresolved at 0.2” resolution, which means that their dust disk radii are 20 au (Ansdell et al., 2018). For a handful of these compact disks, high-resolution images were taken, resolving very small compact disks with a radius of only 3-5 au (van der Marel et al., 2022). This is also seen in compact disks in Taurus (Kurtovic et al., 2021). It is currently unclear whether all unresolved disks are equally compact, and actual sizes may have a range between 2-20 au. However, the dust disk fluxes at 1.3mm in Lupus of these unresolved disks range between 0.4-12 mJy (Ansdell et al., 2018). An extrapolation of the size-luminosity relation in Lupus predicts dust disk radii between 2-10 au for this flux range (Hendler et al., 2020; Tazzari et al., 2021), which is similar to the range predicted by drift-dominated disk models (Rosotti et al., 2019, Appelgren et al., subm.), as demonstrated in Figure 1. Also, dust evolution models predict a decrease in observable dust mass of at least 1 order of magnitude in 1 Myr in disks without pressure bumps (Pinilla et al., 2020a; Appelgren et al., 2023) due to radial drift, i.e. by a decrease of the dust disk size.
@@ 2.1 Pebble mass inventory in observed compact protoplanetary disks
In this work, we assume that planet formation takes place in compact dust disks of 20 au radius. This means that we have to review carefully what is known about the dust masses in these disks to be able to use a representative pebble mass content for our pebble accretion model. In most observational disk studies, the dust mass is derived directly from the integrated millimeter flux at 1 mm, using ALMA observations (e.g. Ansdell et al., 2018; Manara et al., 2023). The dust mass is derived under the assumption of optically thin emission and a constant dust temperature of 20 K, using the following relation:
@@
where the dust grain opacity is 2.3 cm2g-1, Bν(Tdust) is the Planck function for a dust temperature of 20 K and is the distance in pc. This dust mass relies on several assumptions which may not be representative for compact disks, as emission is likely more optically thick and the dust temperature is higher as the dust content may be located much closer to the star (van der Marel & Pinilla, 2023). Therefore, one cannot simply use such dust masses from the literature as actual measurements of the available pebble mass in planet formation models. One of the main problems is that for the majority of these disks, the ALMA images remain spatially unresolved at moderate resolution of 0.2” of disk surveys, so that only an upper limit on the dust disk radius can be derived, which means that optical depth and dust temperature remain uncertain. The handful of compact disks imaged at high angular resolution (0.05”) do reveal dust disk radii as small as 3-5 au (Facchini et al., 2019; Kurtovic et al., 2021; van der Marel et al., 2022). As explained above, we assume that the majority of unresolved dust disks in Lupus are similarly small.
@@
Rather than attempting to derive more accurate dust masses from unresolved disk images, we will use the available integrated millimeter fluxes and the limits on their dust disk size for the compact disks in Lupus around low-mass stars (Ansdell et al., 2018; van der Marel et al., 2022) to compare with the expected millimeter flux of a forward model of the simulated disks in this study, to demonstrate that our pebble mass inventory is representative for observed disks at 1-2 Myr. We emphasize that dust disks in Lupus (which are considered to be 1-2 Myr old) have likely already experienced significant mass loss due to radial drift (Pinilla et al., 2020b; Appelgren et al., 2023) and the initial disk dust masses and dust disk radii right after disk formation are expected to be much larger (Appelgren et al. subm.). Therefore, we assume that the incoming pebble flux originates from a much larger, initial dust disk of 20 au radius, whereas the dust disk of the simulation is only 2-4 au in radius, which is the disk region where the embryos for the planet formation model will be distributed (see Section 3), as a snapshot at 1-2 Myr where pebble drift has already taken place. This means that we forward model the expected fluxes along an evolutionary track of dust disks that are just formed and then quickly decrease in mass and size as the result of radial drift over 1-2 Myr, in two steps.
@@
We utilized a generic protoplanetary disk model in RADMC-3D, configuring the surface density of the dust disk to match the pebble surface density in our model (see Section 2.2). Within the model, we use stellar masses of 0.1 M⊙, 0.3 M⊙, and 0.5 M⊙, accompanied by luminosities of 0.067 L⊙, 0.33 L⊙, and 0.65 L⊙ (Baraffe et al., 2015), and assumed dust outer radii of 2 au, 3 au, and 4 au, respectively for the very compact dust disk of the simulation. These luminosities, in conjunction with effective temperatures derived from Alcalá et al. (2017) corresponding to the stellar masses (3125 K, 3270 K, and 3342 K, respectively), determined both the stellar radius and the inner dust disk radius, assumed to be equal to the dust sublimation radius (T=1500 K). The dust opacities were computed using the optool package (Dominik et al., 2021), employing DSHARP dust particle opacities and setting the minimum particle size to amin as 0.050 m and the maximum size as amax as 3mm (Birnstiel et al., 2018). Anisotropic scattering was assumed within RADMC-3D to enhance the fidelity of reproducing observed emissions from compact disks at 1.3 mm.
The volumetric density distribution in RADMC-3D is defined as follows:
where is the radial distance to the star from the disk, (r) is the dust scale height of the disk, and , is the dust surface density defined as:
where is the power law index and , with and the dust mass and outer radius of the dust disk, respectively. The index is selected to match the model pebble density profile from Section 2.2 in Eq. 4. The input dust mass is then derived by integrating the profile out to an outer dust radius of 2, 3 or 4 au as explained above for stellar masses of 0.1, 0.3, and 0.5 M⊙, resulting in dust masses of 1.5, 10, and 12 M⊕, respectively. For such compact disks, at the end of the radial drift process, the index . To derive the expected millimeter flux, we set a distance of 160 parsecs, similar to the Lupus star forming region.
@@
Figure 2 shows all the available fluxes from the observed protoplanetary disks in Lupus (Ansdell et al., 2018), with stellar masses less than 0.5 M⊙, for direct comparison with our models. As explained above, for the majority of these disks, the size can only be constrained to be 20 au and therefore represent a range of dust disk sizes and dust disk masses along the evolutionary trajectory of radial drift, considering their range of fluxes and the size-luminosity relation (Figure 1).
It is evident from Figure 2 that both disks with the initial pebble reservoir and the ones with a more advanced dust evolutionary stage fall within the typical flux range observed in Lupus disk observations. In the later evolutionary stage, the most outer planet may already be formed, but at orbital radii in the most inner part of the disk well within 1 au. During the simulation, the embryos in this region of the disk will continue to accrete material until reaching the pebble isolation mass, at which point radial drift is halted outside the most exterior planet. This process may result in a very compact disk 3-5 au with one or more pressure bumps depending on the reached planet masses. Such a dust disk would have a significantly lower dust mass than the initial disk of 20 au, but both masses still fall within the range of unresolved disks seen in Lupus (Figure 2). These compact disks possibly constitute the majority of disk observations within the submillimeter range for stellar masses M⊙.

### v3

@@ Abstract
Context. Due to the improvements in radial velocity and transit techniques, we know that rocky or rocky-icy planets, in particular close-in super-Earths in compact configurations, are the most common ones around M dwarfs. On the other hand, thanks to the high angular resolution of ALMA we know that many disks around very low-mass stars (between 0.1 and 0.5 M⊙) are rather compact and small (without observable substructures and radius less than 20 au), which favors the idea of an efficient radial drift that could enhance planet formation in the terrestrial zone.
Aims. Our aim was to investigate the potential formation paths of the observed close-in rocky exoplanet population around M dwarfs, especially close-in super-Earths, assuming that planet formation could take place in compact disks with an efficient dust radial drift.
Methods. We developed N-body simulations that include a sample of embryos growing by pebble accretion exposed to planet-disk interactions, star-planet tidal interactions, and general relativistic corrections that include the evolution of the luminosity, radius, and rotational period of the star. For a star of 0.1 M⊙, we considered different gas disk viscosities and initial embryo distributions. We also explored planet formation by pebble accretion around stars of 0.3 M⊙ and 0.5 M⊙. Lastly, for each stellar mass, we ran simulations that include a sample of embryos growing by planetesimal accretion instead of pebble accretion.
Results. Our main result is that the sample of simulated planets that grow by pebble accretion in a gas disk with low viscosity () can reproduce the close-in low-mass exoplanet population around M dwarfs in terms of multiplicity, mass, and semi-major axis. Furthermore, we found that a gas disk with high viscosity (), and thus lower pebble accretion rates, cannot reproduce the observed planet masses as no planet more massive than 0.5 M⊕ could be formed in our simulations. In addition, we show that planetesimal accretion favors the formation of smaller planets than pebble accretion does. Whether this planet population truly exists remains unknown with the current instrumental sensitivity.
Conclusions. Rocky planet formation around M dwarfs can take place in compact and small dust disks driven by an efficient radial drift in a gas disk with low viscosity (). This result points toward a new approach in the direction of the disk conditions needed for rocky planet formation around very low-mass stars.
@@
M dwarfs are the most abundant objects in our solar neighborhood, representing almost 70 of the stars in our Galaxy (Chabrier, 2003). Thanks to the improvements in radial velocity and transit detection techniques, we know that they are hosts to closely packed planetary systems with multiple rocky planets (e.g., Shields et al., 2016; Mulders, 2018). To date, rocky planets have been found to be the most abundant around M dwarfs (Sabotta et al., 2021), frequently found close in to the star (Ment & Charbonneau, 2023). These planets are of great interest as many of them are located in the habitable zone of their systems (e.g., Kasting et al., 1993; Kopparapu et al., 2013), which makes M dwarfs ideal targets for the search for life in the Universe.
@@
Even though several authors have addressed the study of Earth-like planets around M dwarfs with both population synthesis models and N-body simulations (e.g., Dugaro et al., 2016; Coleman et al., 2019; Miguel et al., 2020; Liu et al., 2020; Sánchez et al., 2022), the formation of super-Earths (planets with sizes between Earth and Neptune) around M dwarfs is still under debate; it is challenging to form such planets in situ with the classic planet formation theory and based on the low disk masses typically derived from observations (e.g., Raymond et al., 2007; Ciesla et al., 2015; Schlichting, 2018). Thus, it is necessary to address the problem taking into account different premises. For example, Liu et al. (2019) developed a pebble-driven core accretion model in a population synthesis code to study the formation and evolution of planets around stars from the substellar mass limit at 0.08 M⊙ up to 1 M⊙. They found that the characteristic mass of a super-Earth is set by the pebble isolation mass, and that it increases linearly with the mass of the stellar host, and that planets with masses lower than 20 M⊕ can be formed around stars with a wide range of metallicities.However, they did not consider gravitational interactions among the sample of embryos, which is of great importance as collisions among embryos play a primary role in defining the final masses and orbital configurations of planetary systems (e.g., Chambers, 2001; O’Brien et al., 2006; Raymond et al., 2007; de Elía et al., 2013; Ronco & de Elía, 2018; Sánchez et al., 2022). Thus, their simulations may not be realistic when compared to the final exoplanet properties. Other studies, such as Mulders et al. (2021), proposed a pebble-drift and accretion model to explain the occurrence rate of transiting planets around M dwarfs. They found that the fraction of close-in super-Earths is higher around lower-mass stars and matches the exoplanet occurrence rates from Kepler. However, their results are based on the study of a two-planet system, and they do not take into account planet-disk interactions. The inclusion of gas disk interactions with a sample of embryos is crucial, as is defining the early dynamical evolution and orbital configuration of planetary systems (e.g., Tanaka & Ward, 2004; Paardekooper et al., 2011; Ida et al., 2020). Moreover, the inclusion of star-planet tidal interactions is also relevant in the long-term dynamical evolution of close-in planets around very low-mass stars due to the fast evolution of the stellar radius and rotation period (e.g., Bolmont et al., 2011; Sánchez et al., 2020).
@@
Generally, the assumptions regarding the disk properties in planet population models may not be representative when compared with the latest insights from disk observations. Recently, the Atacama Large
@@
Millimeter/sub-millimeter Array (ALMA) disk surveys have shown the existence of compact dust disks around very low-mass stars (e.g., Facchini et al., 2019; Kurtovic et al., 2021; van der Marel et al., 2022). A handful of disks are truly as small as 3-5 au in radius, whereas many of these disks remain unresolved in current ALMA images and their radius is only constrained to be 20 au. Such compact disks, lacking large-scale substructures, are more often found around low-mass stars than high-mass stars (van der Marel & Mulders, 2021) and are expected to be dominated by radial drift and thus a high pebble flux (Pinilla et al., 2013; Banzatti et al., 2020, 2023). In such disks, super-Earths are expected to form, consistent with a formation through pebble accretion in drift-dominated disks (Mulders et al., 2021; van der Marel & Mulders, 2021). Recently, Chachan & Lee (2023) used a simple framework that combines the theory of pebble accretion with the measurements of dust masses in protoplanetary disks around late M dwarfs. They demonstrated that the
@@
FGK dwarfs (e.g., Dressing et al., 2015; Mulders et al., 2015), in agreement with van der Marel & Mulders (2021). Additionally, by modeling the evolution of a stellar cluster, Appelgren et al. (2023) investigated the efficiency of radial drift in clearing the disk of millimeter-sized particles. Throughout the disk’s lifetime, they found an agreement between the mass locked in pebbles and the evolution of dust masses and ages inferred from nearby star-forming regions. They stated that disks are born with dust masses of up to around 50 M⊕ for low-mass stars of 0.1 M⊙, and quickly decreases to 1 M⊕ at 1-2 Myr due to radial drift, which suggests that at the earliest stages there could be enough material to form inner planetary cores. Such high disk dust masses have indeed been inferred for embedded disks in the Class I stage (Tychoniec et al., 2020).
Motivated by the fact that the formation of planets more massive than the Earth is hard to explain with classical planet formation models, and that dust disks around very low-mass stars are rather small and compact, dominated by radial drift and with an initial dust mass high enough to form inner planetary cores, we investigated different rocky planet formation paths around M dwarfs. In order to do that, we assumed that planet formation takes place in compact dust disks, dominated by radial drift rather than the well-studied gapped disks (e.g., Andrews et al., 2018) where dust traps are thought to halt the radial drift (Pinilla et al., 2012). Based on this assumption, we developed different sets of N-body simulations that include a sample of small embryos growing by pebble accretion that experience planet-disk interactions as well as star-planet tidal interactions that incorporate the evolution of the luminosity, radius and rotational period of the star. All simulations run for 50 Myr to take into account late collisions among embryos, considering a gas disk lifetime of 10 Myr. We explored different planet formation scenarios around stars of 0.1 M⊙, 0.3 M⊙ and 0.5 M⊙.
This work is structured as follows. In Section 2 we describe the dust disk model used based on compact dust disk observations. In Section 3 we describe the planet formation model included in the N-body code, and present different formation scenarios. Appendix A and B summarize the details of the gas disk evolution model and external forces on the N-body code. In Section 4 we show the resulting simulated planetary systems and explain their orbital history and final configurations. In Section 5 we compare the simulated planets with the confirmed close-in low-mass exoplanet population around M dwarfs. Finally, in Section 6 we summarize some points of discussion and in Section 7 we give the main conclusions of our work.
@@ 2 Compact dust disks: Models and observations
Current ALMA observations of protoplanetary disks around M dwarfs at 1-2 Myr primarily show compact dust continuum disks that remain unresolved at 0.2” resolution, which means that their dust disk radii are 20 au (Ansdell et al., 2018). For a handful of these compact disks, high-resolution images were taken, resolving very small compact disks with a radius of only 3-5 au (van der Marel et al., 2022). This is also seen in compact disks in Taurus (Kurtovic et al., 2021). It is currently unclear whether all unresolved disks are equally compact, and actual sizes may have a range between 2-20 au. However, the dust disk fluxes at 1.3mm in Lupus of these unresolved disks range between 0.4-12 mJy (Ansdell et al., 2018). An extrapolation of the size-luminosity relation in Lupus predicts dust disk radii between 2-10 au for this flux range (Hendler et al., 2020; Tazzari et al., 2021), which is similar to the range predicted by drift-dominated disk models (Rosotti et al., 2019, Appelgren et al., subm.), as demonstrated in Figure 1. Dust evolution models also predict a decrease in observable dust mass of at least one order of magnitude in 1 Myr in disks without pressure bumps (Pinilla et al., 2020a; Appelgren et al., 2023) due to radial drift (i.e., by a decrease in the dust disk size).
@@ 2.1 Pebble mass inventory in observed compact protoplanetary disks
In this work, we assume that planet formation takes place in compact dust disks of 20 au radius. This means that we have to review carefully what is known about the dust masses in these disks to be able to use a representative pebble mass content for our pebble accretion model. In most observational disk studies, the dust mass is derived directly from the integrated millimeter flux at 1 mm, using ALMA observations (e.g., Ansdell et al., 2018; Manara et al., 2023). The dust mass is derived under the assumption of optically thin emission and a constant dust temperature of 20 K, using the following relation:
@@
Here the dust grain opacity is 2.3 cm2g-1, Bν(Tdust) is the Planck function for a dust temperature of 20 K, and is the distance in pc. This dust mass relies on several assumptions which may not be representative for compact disks, as emission is likely more optically thick and the dust temperature is higher as the dust content may be located much closer to the star (van der Marel & Pinilla, 2023). Therefore, one cannot simply use such dust masses from the literature as actual measurements of the available pebble mass in planet formation models. One of the main problems is that for the majority of these disks, the ALMA images remain spatially unresolved at moderate resolution of 0.2” of disk surveys, so that only an upper limit on the dust disk radius can be derived, which means that optical depth and dust temperature remain uncertain. The handful of compact disks imaged at high angular resolution (0.05”) do reveal dust disk radii as small as 3-5 au (Facchini et al., 2019; Kurtovic et al., 2021; van der Marel et al., 2022). As explained above, we assume that the majority of unresolved dust disks in Lupus are similarly small.
@@
Rather than attempting to derive more accurate dust masses from unresolved disk images, we use the available integrated millimeter fluxes and the limits on their dust disk size for the compact disks in Lupus around low-mass stars (Ansdell et al., 2018; van der Marel et al., 2022) to compare with the expected millimeter flux of a forward model of the simulated disks in this study, to demonstrate that our pebble mass inventory is representative for observed disks at 1-2 Myr. We emphasize that dust disks in Lupus (which are considered to be 1-2 Myr old) have likely already experienced significant mass loss due to radial drift (Pinilla et al., 2020b; Appelgren et al., 2023) and the initial disk dust masses and dust disk radii right after disk formation are expected to be much larger (Appelgren et al., 2024). Therefore, we assume that the incoming pebble flux originates from a much larger, initial dust disk of 20 au radius, whereas the dust disk of the simulation is only 2-4 au in radius, which is the disk region where the embryos for the planet formation model will be distributed (see Section 3), as a snapshot at 1-2 Myr where pebble drift has already taken place. This means that we forward model the expected fluxes along an evolutionary track of dust disks that are just formed and then quickly decrease in mass and size as the result of radial drift over 1-2 Myr, in two steps.
@@
We utilized a generic protoplanetary disk model in RADMC-3D, configuring the surface density of the dust disk to match the pebble surface density in our model (see Section 2.2). Within the model, we used stellar masses of 0.1 M⊙, 0.3 M⊙, and 0.5 M⊙, accompanied by luminosities of 0.067 L⊙, 0.33 L⊙, and 0.65 L⊙ (Baraffe et al., 2015), and assumed dust outer radii of 2 au, 3 au, and 4 au, respectively for the very compact dust disk of the simulation. These luminosities, in conjunction with effective temperatures derived from Alcalá et al. (2017) corresponding to the stellar masses (3125 K, 3270 K, and 3342 K, respectively), determined both the stellar radius and the inner dust disk radius, assumed to be equal to the dust sublimation radius (T=1500 K). The dust opacities were computed using the optool package (Dominik et al., 2021), employing DSHARP dust particle opacities and setting the minimum particle size to amin as 0.050 m and the maximum size as amax as 3 mm (Birnstiel et al., 2018). Anisotropic scattering was assumed within RADMC-3D to enhance the fidelity of reproducing observed emissions from compact disks at 1.3 mm.
The volumetric density distribution in RADMC-3D is defined as:
where is the radial distance to the star from the disk, (r) is the dust scale height of the disk, and is the dust surface density, defined as
where is the power law index and , with and the dust mass and outer radius of the dust disk, respectively. The index is selected to match the model pebble density profile from Section 2.2 in Eq. 4. The input dust mass is then derived by integrating the profile out to an outer dust radius of 2, 3 or 4 au as explained above for stellar masses of 0.1, 0.3, and 0.5 M⊙, resulting in dust masses of 1.5, 10, and 12 M⊕, respectively. For such compact disks, at the end of the radial drift process, the index . To derive the expected millimeter flux, we set a distance of 160 parsecs, similar to the Lupus star-forming region.
@@
Figure 2 shows all the available fluxes from the observed protoplanetary disks in Lupus (Ansdell et al., 2018), with stellar masses less than 0.5 M⊙, for direct comparison with our models. As explained above, for the majority of these disks, the size can only be constrained to be 20 au, and therefore represent a range of dust disk sizes, and dust disk masses along the evolutionary trajectory of radial drift, considering their range of fluxes, and the size-luminosity relation (Figure 1).
It is evident from Figure 2 that both disks with the initial pebble reservoir, and the ones with a more advanced dust evolutionary stage, fall within the typical flux range observed in Lupus disk observations. In the later evolutionary stage, the most outer planet may already be formed, but at orbital radii in the most inner part of the disk well within 1 au. During the simulation, the embryos in this region of the disk will continue to accrete material until reaching the pebble isolation mass, at which point radial drift is halted outside the most exterior planet. This process may result in a very compact disk 3-5 au with one or more pressure bumps depending on the reached planet masses. Such a dust disk would have a significantly lower dust mass than the initial disk of 20 au, but both masses still fall within the range of unresolved disks seen in Lupus (Figure 2). These compact disks possibly constitute the majority of disk observations within the submillimeter range for stellar masses M⊙.